OpenStudy (anonymous):

Solve this system of equations: 2x - y = 7 5x + y = 21 A. (4,1) B. (5,3) C. (3,6) D. (2,-3)

OpenStudy (uri):

2x - y = 7 5x + y = 21 ========= 7x=28 x=28/7 x=4 5(4)+y=21 20+y=21 y=21-20 y=1 Answer is (4,1) (A)

OpenStudy (anonymous):

can toy explain this one too Solve by substitution method: y = -2x + 3 3x - 2y = 8 A. no solution B. (14, -25) C. (-14, 31) D. (2, -1) @uri

OpenStudy (anonymous):

D

OpenStudy (anonymous):

same as upper one